OpenStudy (anonymous):

using complete sentences explain how to find the minimum value for each for each function and and determine which function has the smallest y value

OpenStudy (anonymous):

f(x) = 3x^2 + 12x + 16 And g(x) = 2sin (2x-pi) +4

OpenStudy (anonymous):

@ejones3178 @amistre64 @wyattp17

OpenStudy (anonymous):

@acxbox22

OpenStudy (anonymous):

@tHe_FiZiCx99 @geerky42

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

dont u graph it and see the amplitude and midline

OpenStudy (loser66):

ok, may be you are not cal student. I make another question: for f(x), it is an upward parabola, right? so, its min is its vertex, what is it? ( I mean coordinate of the vertex)

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

not suree

OpenStudy (loser66):

hmmmmm means "don't know", right?

OpenStudy (loser66):

lol, x-coordinate is -b/2a =??

OpenStudy (anonymous):

Yess

OpenStudy (loser66):

=? , number, please

OpenStudy (anonymous):

oh sorry haha ok so

OpenStudy (anonymous):

wow i havent done that in a while.. Is it 16

OpenStudy (loser66):

nope, ax^2 +bx+c , your case 3x^2+12x +16, so a =3, b =12 c =16 ( but we don't use c, let it aside) so that what is -b/2a?

OpenStudy (anonymous):

-12/6

OpenStudy (loser66):

=??

OpenStudy (anonymous):

2

OpenStudy (loser66):

nope, you forgot the - sign, redo

OpenStudy (anonymous):

-2?

OpenStudy (loser66):

ok, now we have x-coordinate of the vertex, plug x =-2 into f(x) , what do you get?

OpenStudy (anonymous):

-4

OpenStudy (loser66):

it's not correct, show me your work so that I can see what 's wrong

OpenStudy (anonymous):

3(-2)^2 +12(-2) + 16

OpenStudy (anonymous):

-6^2 -24 + 16

OpenStudy (anonymous):

36-24+16

OpenStudy (loser66):

hihihih 3(-2)^2 = 3*4 =12, it's not -6^2. You must square the number inside the bracket first, then time 3 from the out side

OpenStudy (loser66):

so, it is 12 -24+16 =-4

OpenStudy (loser66):

So far, from f(x) you have the vertex (-2,-4) right?

OpenStudy (loser66):